# Distilling Textual Priors from LLM to Efficient Image Fusion

### v2

@@ Abstract
Multi-modality image fusion aims to synthesize a single, comprehensive image from multiple source inputs. Traditional approaches, such as CNNs and GANs, offer efficiency but struggle to handle low-quality or complex inputs. Recent advances in text-guided methods leverage large model priors to overcome these limitations, but at the cost of significant computational overhead, both in memory and inference time. To address this challenge, we propose a novel framework for distilling large model priors, eliminating the need for text guidance during inference while dramatically reducing model size. Our framework utilizes a teacher-student architecture, where the teacher network incorporates large model priors and transfers this knowledge to a smaller student network via a tailored distillation process. Additionally, we introduce spatial-channel cross-fusion module to enhance the model’s ability to leverage textual priors across both spatial and channel dimensions. Our method achieves a favorable trade-off between computational efficiency and fusion quality. The distilled network, requiring only 10% of the parameters and inference time of the teacher network, retains 90% of its performance and outperforms existing SOTA methods. Extensive experiments demonstrate the effectiveness of our approach. Codes are available at https://github.com/Zirconium233/DTPF
Keywords Image Fusion  $\cdot$ Knowledge Distillation  $\cdot$ Large Language Models  $\cdot$ Multi-modality Learning
1 Introduction
Image fusion plays a important role in visual enhancement within digital image processing. For example, visible-infrared image fusion combines color-based details from visible images, which are easily interpretable by humans, with radiation-based features from infrared images, which are highly effective for target detection and operations under low-light conditions. By integrating complementary information from both modalities, this fusion method produces high-quality images that enhance both human visual interpretation and machine-based detection performance.
@@ 1 Introduction
During the imaging process, environmental and device-related constraints often degrade the quality of source images. Visible images may suffer from low resolution, while infrared images are prone to various noise types. Traditional fusion methods, including FusionGAN [1], U2Fusion [2], and SwinFusion [3], fuse source images without effectively distinguishing between degraded information and meaningful image content, leading to suboptimal performance. Some approaches [4] mitigate these issues by relying on manual pre-processing to enhance source images, but their flexibility is limited. Recently, text-guided methods, such as TextFusion and Text-IF [5], have emerged as a promising solution to address these challenges. These methods leverage the capabilities of multimodal large language models (MLLMs) to generate text captions or degradation descriptions (e.g., low light) for source images, enabling the model to adaptively distinguish image content from degradation and enhance image quality. By integrating these priors, the model gains enhanced semantic understanding of the image content, which effectively promote the fusion processes. For instance, Text-IF utilizes MLLMs [6] to generate task lists and incorporates CLIP [7] to guide the fusion process based on these descriptions. Similarly, TextFusion [8] employs text-based interactions to build models capable of focusing specific image elements, such as emphasizing trees over people, based on captions generated by LLMs or human input.
@@
In response to these issues, we propose a large-model prior distillation framework that eliminates the need for text guidance during inference while significantly reducing model size, as shown in Figure 1. Our framework employs a teacher-student architecture, where the teacher network integrates MLLM priors via a feature modulation operations. To fully exploit textual priors across both spatial and channel dimensions, we introduce spatial-channel cross-fusion module (SCFM). A specially tailored prior distillation loss is then used to transfer textual knowledge from both feature and output perspectives, enabling the student network to mimic the teacher’s feature processing without requiring textual inputs. By leveraging this approach, our method achieves a 90% reduction in model size (excluding CLIP, which alone requires five times the parameters of the original fusion module) while retaining 90% of the teacher model’s performance.
@@
  • We introduce spatial-channel cross fusion module to fully exploit textual information across both spatial and channel dimensions, improving the model’s ability to leverage textual priors.
@@ 3.1 Framework
However, the computational cost of LLMs and CLIP is substantial. To mitigate this, we propose a prior distillation framework consisting of a text-guided teacher network $\Phi_{t}$ and its distilled student network $\Phi_{s}$. During training, $\Phi_{t}$ utilizes textual information from MLLMs, which is then transferred to $\Phi_{s}$ through our tailored prior distillation loss. Since semantic information is encoded across both spatial and channel dimensions, we introduce spatial-channel cross fusion blocks within the network. As illustrated in Figure 2, this framework enables the student network to internalize textual information while significantly reducing model size. Both networks share similar structural blocks, with the student network being more compact.
3.2 Network Architecture
@@
3.2.1 Text-Guided Teacher Network
The teacher network adopts a hierarchical structure with four levels of feature processing. Given its critical role in leveraging textual priors, the core operator aims to effectively modeling feature in spatial and channel dimensions [16]. Additionally, we introduce spatial-channel cross fusion blocks as the core operator, enabling the exploration of correlations between spatial and channel dimensions. Text information is fused through the sequential feature modulation, ensuring effective integration of textual priors. This network is organized into three primary components:
1) Dual-Stream Encoder: Dual-Stream Encoder: Processes visible and infrared inputs through parallel branches, each consisting of overlapped patch embedding layers and spatial and channel self-attention blocks. The hierarchical feature transformation at level $l$ is expressed as:
@@ 3.2.1 Text-Guided Teacher Network
where $\mathcal{F}_{k}$ represents the $k$-th spatial-channel self-attention block for modality $i\in{vis,ir}$, and $\circ$ denotes function composition. Each block integrates sequential Windowed Self-Attention Blocks (SSAB) and Transposed Self-Attention Blocks (TSAB) for comprehensive feature encoding:
@@
$\mathbf{F}^{\hat{vis}}_{l}=\text{Conv}{1\times 1}\circ C(\text{SSAB}(\mathbf{F}^{vis}_{l}),\text{TSAB}(\mathbf{F}^{ir}_{l})),$ (6)
$\mathbf{F}^{\hat{ir}}_{l}=\text{Conv}{1\times 1}\circ C(\text{SSAB}(\mathbf{F}^{ir}_{l}),\text{TSAB}(\mathbf{F}^{vis}_{l})),$ (7)
@@
$\mathbf{F}_{out}=\mathcal{R}_{m}\circ\mathcal{R}_{m-1}\circ...\circ\mathcal{R}_{1}(\mathbf{F}_{dec}^{1})$ (11)
Each block $\mathcal{R}_{k}$ consists of SSAB and TSAB operations shared block number with the encoder at corresponding level. At the end, the $\mathbf{F}_{out}$ is the finial result $\mathbf{I}_{fused}$.
3.2.2 Lightweight Student Network
@@
1) Dimension reduction in transformer blocks (from 48 to 16 channels) 2) Removal of text guidance module, directly using $\mathbf{F}_{fused}^{l}$ instead of $\hat{\mathbf{F}_{fused}^{l}}$
3.3 Prior Knowledge Distillation
@@
3.3.1 Teacher Network Training
@@
$\mathcal{L}_{tea}=\lambda_{s}^{t}\mathcal{L}_{ssim}+\lambda_{i}^{t}\mathcal{L}_{int}+\lambda_{g}^{t}\mathcal{L}_{grad}+\lambda_{c}^{t}\mathcal{L}_{color}$ (12)
where $\mathcal{L}_{ssim}$ measures structural similarity through SSIM index, $\mathcal{L}_{int}$ preserves maximum intensity information between source images, $\mathcal{L}_{grad}$ ensures gradient consistency using Sobel operators, and $\mathcal{L}_{color}$ maintains color fidelity in YCbCr space. The weights $\lambda^{t}$ are dynamically adjusted based on text guidance. The details of these losses can be found in the supplementary material.
3.3.2 Progressive Knowledge Transfer
@@
Table 1: Quantitative comparison with SOTA methods on IVF tasks. The number in () is non-trainable parameters but used in inference period. Bold and underlined values indicate the best and second-best results respectively. Params are measured in millions (M).
Method Params(M) MSRS M3FD RoadScene
[…]
SuperFusion 11.23 6.587 0.813 0.557 6.726 0.664 0.522 6.990 0.608 0.452
CDDFuse 2.40 6.701 0.819 0.548 7.070 0.802 0.613 7.475 0.610 0.450
IFCNN 0.08 5.975 0.579 0.479 6.935 0.685 0.590 7.222 0.591 0.536
U2Fusion 0.63 5.246 0.512 0.391 6.872 0.673 0.578 6.739 0.564 0.506
SwinFusion 13.04 6.619 0.825 0.558 6.844 0.746 0.616 7.000 0.614 0.450
PSLPT 3.06 6.307 0.753 0.553 7.204 0.958 0.321 7.077 0.134 0.171
TC-MOA 7.24(+329) 6.633 0.811 0.565 6.747 0.579 0.508 7.387 0.577 0.477
Text-IF 63.8(+151) 6.729 1.051 0.690 6.849 0.780 0.550 7.332 0.739 0.578
Ours-teacher 40.3(+151) 6.749 1.060 0.732 6.965 0.896 0.706 7.248 0.743 0.639
Ours-distilled 4.10 6.763 1.075 0.734 6.989 0.927 0.704 7.279 0.751 0.634
4 Experiments
@@
4.1 Implementation Details and Datasets
Implementation Details: Both teacher and student networks are trained using AdamW optimizer with a learning rate of 0.0001. Input images are cropped to $128\times 128$ patches during training. For the teacher network without text guidance, the hyper-parameters $\{\lambda_{i},\lambda_{s},\lambda_{g},\lambda_{c}\}$ are set to $\{24,40,48,12\}$ respectively. When text guidance is enabled, task-specific parameters are provided in the supplementary material. We employ QWen2VL [19], a open source large vision-language model, to generate descriptive text labels for input image pairs. All experiments are conducted on four NVIDIA GeForce RTX 4090 GPUs using PyTorch framework.
@@ 4.2 Comparison on IVF Tasks
Table 1 presents the quantitative comparison with SOTA methods on three IVF datasets. Our teacher network consistently outperforms existing methods across all metrics and datasets, demonstrating the effectiveness of text-guided fusion. On the MSRS dataset, our teacher network achieves significant improvements in $Q^{AB/F}$ compared to the previous best method Text-IF.
Figure 3: Compare with different methods on FLIR_05767.jpg in RoadScene datasets. More can be found in appendix.
@@
More remarkably, our distilled student network not only maintains but sometimes exceeds the teacher’s performance, particularly in EN and $Q^{AB/F}$ metrics, while reducing parameters by 90%. This suggests that the knowledge distillation process effectively transfers the teacher’s knowledge to a much smaller model. The slight performance variations between teacher and student networks can be attributed to the extended training during the distillation process, where the student benefits from both the teacher’s guidance and task-specific objectives. Another reason is that current domain models are generally over-parameterized compared to relatively small datasets, which is detailed in the ablation experiments. Figure 3 shows qualitative comparisons on challenging cases from different datasets. Our method better preserves thermal information from infrared images while maintaining visible details and natural appearance, especially in scenarios with extreme lighting conditions or complex textures.

### v3

@@ Abstract
Multi-modality image fusion aims to synthesize a single, comprehensive image from multiple source inputs. Traditional approaches, such as CNNs and GANs, offer efficiency but struggle to handle low-quality or complex inputs. Recent advances in text-guided methods leverage large model priors to overcome these limitations, but at the cost of significant computational overhead, both in memory and inference time. To address this challenge, we propose a novel framework for distilling large model priors, eliminating the need for text guidance during inference while dramatically reducing model size. Our framework utilizes a teacher-student architecture, where the teacher network incorporates large model priors and transfers this knowledge to a smaller student network via a tailored distillation process. Additionally, we introduce a spatial-channel cross-fusion module to enhance the model’s ability to leverage textual priors across both spatial and channel dimensions. Our method achieves a favorable trade-off between computational efficiency and fusion quality. The distilled network, requiring only 10% of the parameters and inference time of the teacher network, retains 90% of its performance and outperforms existing SOTA methods. Extensive experiments demonstrate the effectiveness of our approach. Codes are available at https://github.com/Zirconium233/DTPF
Index Terms:
Image Fusion, Knowledge Distillation, Large Language Models, Multi-modality Learning.
I Introduction
Image fusion plays an important role in visual enhancement within digital image processing. For example, visible-infrared image fusion combines color-based details from visible images, which are easily interpretable by humans, with radiation-based features from infrared images, which are highly effective for target detection and operations under low-light conditions. By integrating complementary information from both modalities, this fusion method produces high-quality images that enhance both human visual interpretation and machine-based detection performance. This work also shares similarities with pan-sharpening, a multi-modal image fusion task that leverages frequency domain analysis, as explored in recent studies such as [1].
@@ I Introduction
During the imaging process, environmental and device-related constraints often degrade the quality of source images. Visible images may suffer from low resolution, while infrared images are prone to various noise types. Traditional fusion methods, including FusionGAN [2], U2Fusion [3], and SwinFusion [4], fuse source images without effectively distinguishing between degraded information and meaningful image content, leading to suboptimal performance. Some approaches [5] mitigate these issues by relying on manual pre-processing to enhance source images, but their flexibility is limited. Recently, text-guided methods, such as TextFusion [6] and Text-IF [7], have emerged as a promising solution to address these challenges. These methods leverage the capabilities of multimodal large language models (MLLMs) to generate text captions or degradation descriptions (e.g., low light) for source images, enabling the model to adaptively distinguish image content from degradation and enhance image quality. By integrating these priors, the model gains enhanced semantic understanding of the image content, which effectively promotes the fusion processes. For instance, Text-IF utilizes MLLMs [8] to generate task lists and incorporates CLIP [9] to guide the fusion process based on these descriptions. Similarly, TextFusion [6] employs text-based interactions to build models capable of focusing on specific image elements, such as emphasizing trees over people, based on captions generated by LLMs or human input.
@@
In response to these issues, we propose a large-model prior distillation framework that eliminates the need for text guidance during inference while significantly reducing model size, as shown in Fig. 1. Our framework employs a teacher-student architecture, where the teacher network integrates MLLM priors via feature modulation operations. To fully exploit textual priors across both spatial and channel dimensions, we introduce a spatial-channel cross-fusion module (SCFM). A specially tailored prior distillation loss is then used to transfer textual knowledge from both feature and output perspectives, enabling the student network to mimic the teacher’s feature processing without requiring textual inputs. By leveraging this approach, our method achieves a 90% reduction in model size (excluding CLIP, which alone requires five times the parameters of the original fusion module) while retaining 90% of the teacher model’s performance.
@@
  • We introduce a spatial-channel cross-fusion module to fully exploit textual information across both spatial and channel dimensions, improving the model’s ability to leverage textual priors.
@@ III-A Framework
However, the computational cost of LLMs and CLIP is substantial. To mitigate this, we propose a prior distillation framework consisting of a text-guided teacher network $\Phi_{t}$ and its distilled student network $\Phi_{s}$. During training, $\Phi_{t}$ utilizes textual information from MLLMs, which is then transferred to $\Phi_{s}$ through our tailored prior distillation loss. Since semantic information is encoded across both spatial and channel dimensions, we introduce spatial-channel cross fusion blocks within the network. As illustrated in Fig. 2, this framework enables the student network to internalize textual information while significantly reducing model size. Both networks share similar structural blocks, with the student network being more compact.
III-B Network Architecture
@@
III-B1 Text-Guided Teacher Network
The teacher network adopts a hierarchical structure with four levels of feature processing. Given its critical role in leveraging textual priors, the core operator aims to effectively model features in spatial and channel dimensions [17]. Additionally, we introduce spatial-channel cross fusion blocks as the core operator, enabling the exploration of correlations between spatial and channel dimensions. Text information is fused through the sequential feature modulation, ensuring effective integration of textual priors. This network is organized into three primary components:
1) Dual-Stream Encoder: Processes visible and infrared inputs through parallel branches, each consisting of overlapped patch embedding layers and spatial and channel self-attention blocks. The hierarchical feature transformation at level $l$ is expressed as:
@@ III-B1 Text-Guided Teacher Network
[…]
@@
$\mathbf{F}^{\hat{vis}}_{l}=\text{Conv}_{1\times 1}\circ C(\text{SSAB}(\mathbf{F}^{vis}_{l}),\text{TSAB}(\mathbf{F}^{ir}_{l})),$ (6)
$\mathbf{F}^{\hat{ir}}_{l}=\text{Conv}_{1\times 1}\circ C(\text{SSAB}(\mathbf{F}^{ir}_{l}),\text{TSAB}(\mathbf{F}^{vis}_{l})),$ (7)
@@
$\mathbf{F}_{out}=\mathcal{R}_{m}\circ\mathcal{R}_{m-1}\circ\ldots\circ\mathcal{R}_{1}(\mathbf{F}_{dec}^{1})$ (11)
Each block $\mathcal{R}_{k}$ consists of SSAB and TSAB operations sharing block number with the encoder at the corresponding level. At the end, the $\mathbf{F}_{out}$ is the final result $\mathbf{I}_{fused}$.
III-B2 Lightweight Student Network
@@
  • Dimension reduction in transformer blocks (from 48 to 16 channels)
  • Removal of text guidance module, directly using $\mathbf{F}_{fused}^{l}$ instead of $\hat{\mathbf{F}_{fused}^{l}}$
III-C Prior Knowledge Distillation
@@
III-C1 Teacher Network Training
@@
$\mathcal{L}_{int}=\frac{1}{HW}\|I_{f}-\max(I_{vis}^{g},I_{ir}^{g})\|_{1}$ (13)
where $F_{CbCr}$ denotes the transfer function of RGB to YCbCr.
These loss terms work together to ensure high-quality image fusion: $\mathcal{L}_{int}$ maintains the maximum intensity information from both source images, $\mathcal{L}_{SSIM}$ preserves structural details through the SSIM [20] metric, $\mathcal{L}_{grad}$ ensures edge consistency through gradient preservation, and $\mathcal{L}_{color}$ maintains natural color appearance. The text-dependent weights $\lambda^{t}$ allow dynamic adjustment of each loss component’s contribution based on the specific fusion requirements described in the text prompt.
III-C2 Progressive Knowledge Transfer
@@
IV Experiments
@@
IV-A Implementation Details and Datasets
Implementation Details: Both teacher and student networks are trained using AdamW optimizer with a learning rate of 0.0001. Input images are cropped to $128\times 128$ patches during training. For the teacher network, the hyper-parameters $\{\lambda_{int},\lambda_{ssim},\lambda_{grad},\lambda_{color}\}$ are set to $\{24,40,48,12\}$ respectively. When text guidance is enabled, task-specific parameters are provided in the supplementary material. We employ QWen2VL [21], an open-source large vision-language model, to generate descriptive text labels for input image pairs. All experiments are conducted on four NVIDIA GeForce RTX 4090 GPUs using PyTorch framework.
@@ IV-B Comparison on IVF Tasks
Table I presents the quantitative comparison with SOTA methods on three IVF datasets. Our teacher network consistently outperforms existing methods across all metrics and datasets, demonstrating the effectiveness of text-guided fusion. On the MSRS dataset, our teacher network achieves significant improvements in $Q^{AB/F}$ compared to the previous best method Text-IF.
Figure 3: Qualitative comparison of different image fusion methods on a challenging scene (FLIR_05767.jpg) from the RoadScene dataset. Our method better preserves thermal information from infrared images while maintaining visible details and natural appearance, especially in scenarios with extreme lighting conditions or complex textures. For a more exhaustive visual comparison across various scenarios and methods, please refer to Fig. 8.
@@
More remarkably, our distilled student network not only maintains but sometimes exceeds the teacher’s performance, particularly in EN and $Q^{AB/F}$ metrics, while reducing parameters by 90%. This suggests that the knowledge distillation process effectively transfers the teacher’s knowledge to a much smaller model. The slight performance variations between teacher and student networks can be attributed to the extended training during the distillation process, where the student benefits from both the teacher’s guidance and task-specific objectives. Another reason is that current domain models are generally over-parameterized compared to relatively small datasets, which is detailed in the ablation experiments.